1. The Metro Electric Company produces and distributes electricity to residential customers in New Brunswick.  The demand for electricity is given by

P = 0.04 – 0.01Q

The resulting marginal revenue function is: MR = 0.04 – 0.02Q

Metro’s marginal cost function is: MC = 0.005 + 0.0075Q

where Q = millions of kilowatt hours and P = the dollars/kilowatt hour.

If the mayor of New Brunswick wants to maximize surplus what price per month should she mandate that Metro charge? How many kilowatt hours will be consumed?
Answer #1

Surplus is maximized where firm is attained equilibrium.

Equilibrium condition: MR = MC

0.04 - 0.02Q = 0.005 + 0.0075Q

0.04 - 0.005 = 0.0075Q + 0.02Q

0.035 = 0.0275Q

Q = 0.035/0.0275

Q = 1.27 millions of kilowatt hours

P = 0.04 - 0.01(1.27)

P = 0.04 - 0.0127

P = 0.0273 kilowatt hours